# Perceptions to Beliefs: Exploring Precursory Inferences for Theory of Mind in Large Language Models

### v2

@@ Abstract
Chani Jung  Dongkwan Kim  Jiho Jin  Jiseon Kim Yeon SeonwooYejin Choi♢♠Alice OhHyunwoo Kim KAIST  Amazon  Allen Institute for AI  University of Washington
@@ 1 Introduction
Our contributions are as follows. First, we construct perception-augmented ToM benchmarks which enable the evaluation of the two precursory inferences for ToM in LLMs (§2): perception inference and perception-to-belief inference. Second, using these benchmarks, we show that current LLMs are good at inferring the perceptions of others but struggle to infer beliefs from the perceptual information (§5.1, §5.2, and §5.4). Lastly, we introduce the PercepToM framework to improve LLMs’ ToM reasoning by leveraging their strong perception inference while supplementing their perception-to-belief inference3). Our method enhances LLMs’ performance on ToMi and FANToM, especially on the false belief scenarios (§5.3).
@@ 2.1 Perception Inference and
These are considered as the precursory inferences for ToM (Rakoczy, 2022). We illustrate how they are defined through the Sally-Anne test, a widely used psychology test (Baron-Cohen et al., 1985) for evaluating ToM. In this test’s narrative, Sally does not witness Anne move the marble to the basket, which Sally had previously seen in the box, because Sally left the room. We refer to the capability to infer others’ perceptions (e.g., “Did Sally see the marble being moved to the basket?”) as perception inference. Next, we define the process of deducing others’ beliefs based on their perceptual information (e.g., “Sally did not see the marble move to the basket. Where will Sally look for it when she returns?”) as perception-to-belief inference. However, existing ToM benchmarks mainly focus on the surface-level performance of LLMs on ToM questions. Hence, what is missing in their underlying inference capabilities remains underexplored.
To this end, we construct Percept-ToMi and Percept-FANToM by annotating the character’s perception of each unit of information in the context on top of benchmark ToMi (Le et al., 2019) and FANToM (Kim et al., 2023), respectively. The annotation examples are illustrated in Figure 2.
@@ 2.3 Perception-Augmented ToM Benchmarks
To construct Percept-ToMi, we sample 150 story-question pairs for each of the four ToM question types in ToMi [^1] We use the Fixed and Disambiguated ToMi constructed by Sclar et al. (2023), where sentences are inserted to disambiguate the location of containers in the story, and some mislabeled questions are corrected.: first-order true/false beliefs, and second-order true/false beliefs. We automatically annotate perceivers of the scenes in ToMi using SymbolicToM (Sclar et al., 2023) and manually verify the samples. SymbolicToM tracks the witnesses of each scene by maintaining a graphical representation of the true world state, allowing us to obtain the list of perceivers for each scene from its output. After verifying 50 samples of the automatic annotation, we adjust the perceiver annotations in certain sentence types. Further details are explained in Appendix A.1.
@@
To build Percept-FANToM, we use the entire short conversations in FANToM, but exclude those that cause errors in our perception annotation format. We define the perceivers for each utterance as the people participating in the conversation at that moment (i.e., both speakers and audiences). Based on manual annotations of characters’ joining and leaving times in the conversation by two of the authors, perceivers of each utterance are automatically mapped. Percept-FANToM results in 220 conversations and 735 sets of questions [^2] In FANToM, a set of question asks about the same piece of information within a conversation.. More details are described in Appendix A.2.
@@ 2.4 Task and Evaluation
In order to evaluate the perception inference capability of LLMs, we prompt the models to track characters’ perception of each unit of information in the input context. Specifically, we require the models to respond in the format of a JSON array, which consists of JSON objects containing a unit of information from the context as a key and the perceivers of the information as a value. [^3] We structure the perception inference results in JSON to leverage its parsability and interpretability. Also, recent works use JSON format to improve language model generation quality (Zhou et al., 2023; OpenAI, 2023). We use individual sentences and utterances as the units of information for ToMi and FANToM, respectively. To ensure the generated answers to be in the correct format, we provide an example format of the JSON array using a dummy sentence that does not appear in the datasets. The example input prompt is in Appendix B.1.
To evaluate the model-generated perception inference results, we calculate accuracy for a given input context by determining the ratio of information units for which the model accurately identifies the perceivers. The final perception inference accuracy for a dataset is calculated by taking the average of the accuracies of all contexts in the dataset.
@@
To evaluate the perception-to-belief inference capability of the models, we provide them with a ground truth perception inference result and then query ToM questions from the original benchmarks. The ground truth perception inference result is provided in the same JSON array format we use to evaluate the perception inference capability of LLMs. The example and detailed explanation of the input prompt can be found in Appendix B.2. Since we prompt the model with questions from their original benchmarks, we use the metrics from each original benchmark.
@@ 3 PercepToM: Grounding ToM Reasoning on Perception
According to our experiment results, LLMs perform adequately well in both true and false belief scenarios on perception inference, while they underperform in perception-to-belief inference (§5). Based on these findings, we propose PercepToM, a framework for improving LLM’s ToM reasoning by grounding it in perception information. PecepToM leverages LLM’s strong perception inference capabilities while enhancing their perception-to-belief inference with a simple string-matching rule. PercepToM consists of the following steps as illustrated in Figure 3:
@@
    Perception Inference: The LLM infers which characters perceived each unit of information in the context (e.g., scene description or utterance).
@@
    Perspective Context Extraction: Based on the perception inference result from the LLM, PercepToM extracts the perspective context — i.e., the subset of the input context identified by the LLM as perceived by the target character. This process is conducted by simple string-matching.
@@
If the model correctly performs perception inference, the perspective context will only include what the target character perceived – that is, what they believe to be true, based on the principle of rational belief (Baker et al., 2011). When given this isolated context along with the ToM question, the scenario becomes a simple true belief scenario, wherein the LLM have access to the same information as the target character (i.e., information symmetry).
@@
Figure 3: An overview of our PercepToM framework, which enhances LLMs’ ToM reasoning by: (1) instructing LLMs to infer the perceivers of each information in the context; (2) aiding their perception-to-belief inference through the perspective context extraction step, which isolates the context perceived by the target character; and (3) allowing LLMs to generate responses to ToM questions based on this perspective context.
@@
SymbolicToM (Sclar et al., 2023) also helps LLM’s ToM reasoning by providing only the context included in the target character’s belief state graph to the model. However, constructing the belief graph in SymbolicToM requires manually crafted algorithms tailored to different types of input. In contrast, PercepToM avoids this requirement by leveraging LLM’s perception inference capabilities, which can handle more diverse and complicated contexts, thereby achieving significantly improved generalizability. The example input and output of each step of the algorithm are provided in Appendix C.
4 Experiments
We analyze the perception inference and perception-to-belief inference2.4) performance of LLMs and evaluate our framework on Percept-ToMi and Percept-FANToM.
4.1 Metrics
Perception Inference
In measuring perception inference capability of LLMs, we use the perception inference accuracy introduced in §2.4. In Percept-ToMi, we evaluate the accuracy of all stories, each of which is paired with one ToM question. In FANToM, since a set of questions share the same context, we evaluate models on each context assigned to a set of questions.
Perception-to-Belief Inference and ToM
We evaluate the perception-to-belief inference and ToM performance of LLMs using the original questions and answers from ToMi (Le et al., 2019) and FANToM (Kim et al., 2023). For Percept-ToMi, we measure the accuracy as the ratio of correctly answered questions among all story-question pairs. Note that we do not use the joint accuracy metric proposed in the original ToMi where a story is counted as correctly answered only if all questions about the story are answered correctly. This is because many of the stories in the Fixed and Disambiguated ToMi (Sclar et al., 2023) do not include all six question types of ToMi. For Percept-FANToM, we report the set:ALL score, which requires models to correctly answer all six ToM question types [^4] BeliefQ[Dist.], BeliefQ[Choice], Answerability Q[List], InfoAccess Q[List], Answerability Q[Y/N], InfoAccess Q[Y/N] in each question set.
4.2 Baseline Methods
@@
We compare Vanilla, Chain-of-Thought (CoT; Wei et al., 2022), and System 2 Attention (S2A; Weston and Sukhbaatar, 2023) performance with PercepToM. Vanilla involves LLM directly answering questions based on the given context, while CoT adds the prompt “Let’s think step by step.” to help the model answer ToM questions. S2A improves the reasoning of LLMs by prompting them to extract only the relevant part of the input context before yielding a final response. By using S2A as a baseline, we compare the effectiveness of the perspective context of PercepToM with the relevant context extracted by LLMs using S2A. We also compare with SymbolicToM (Sclar et al., 2023) on ToMi. However, we do not extend this comparison to FANToM, as it is not trivial to apply SymbolicToM to different input formats other than ToMi.
4.3 Target Models
@@
For PercepToM, which leverages the perception reasoning capability of LLMs, we choose models that show reasonable performance on the perception inference task. Specifically, among the eight models, we exclude the bottom two in terms of perception inference accuracy on Percept-FANToM and Percept-ToMi, which are GPT-3.5 Turbo, Claude 3 Haiku, and Gemini 1.0 Pro. As a result, we apply our PercepToM framework to GPT-4 Turbo, GPT-4o, Claude 3 Sonnet, Llama-3 70B Instruct, and Mixtral 8x22B.
5 Results and Discussion
5.1 Perception Inference
As shown in Figure 4, most LLMs exhibit high accuracy on perception inference in both Percept-ToMi and Percept-FANToM. The models’ average perception inference accuracy is 0.781 on Percept-ToMi and 0.926 on Percept-FANToM. Also, they exhibit negligible differences in the accuracy between the true belief and false belief scenarios. In ToMi, all models except for GPT-3.5 Turbo and Gemini 1.0 Pro exhibit a gap of less than 0.1 accuracy between the two scenarios. In FANToM, the accuracy gaps between the two scenarios in all models are no greater than 0.014. This result contrasts with the models’ large performance gap in the two scenarios on ToM questions, suggesting that their limited ToM performance in false belief scenarios is not due to the lack of perception inference capability. Detailed results are in Appendix D.
Especially in the ToMi true belief scenario, the two performances exhibit a near-zero correlation (Figure 5). Although moderate correlations appear in other scenarios, the correlation coefficients are not statistically significant. These results imply that LLMs’ perception inference capability is not directly linked to their ToM performance. This contrasts with human adults, where ToM is strictly dependent on perception inference.
5.2 Perception-to-Belief Inference
LLMs struggle with perception-to-belief inference.
Surprisingly, although the ground truth perception information for all characters are provided in this task, models still underperform in false belief scenarios compared to true belief scenarios (see Figure 4). This trend is consistent with their ToM performance. Moreover, their performances on the perception-to-belief inference task are mostly similar to their ToM performances in all scenarios except for the ToMi true belief scenario. The fact that the LLMs hardly benefit from the additional character perception information, which should serve as significant hints for solving ToM questions, suggests that they have limited capability to infer beliefs from perceptions. The exact performances of models are in Appendix D.
This is consistent across all datasets and scenarios (Figure 5). Notably in FANToM, models exhibit a high correlation between the two performances ($r>0.6$). This correlation likely arises because the two tasks use the same questions. However, since LLMs are showing similar performances in both tasks, we can see that they are not fully leveraging the ground truth perception information in the perception-to-belief inference task.
5.3 PercepToM
@@
Table 1: PercepToM outperforms the baseline models in most of the scenarios on ToMi and FANToM. Bold indicates the best performance within each language model and scenario (true belief or false belief). Performance comparison between PercepToM and SymbolicToM on ToMi can be found in Appendix E.
We also compare the performance of PercepToM and SymbolicToM (Sclar et al., 2023) on ToMi (Appendix E). [^7] Note that SymbolicToM cannot be applied to FANToM as it is tailored to ToMi’s input format. PercepToM performs comparably to SymbolicToM in false belief scenarios across most LLMs. However, in true belief scenarios, SymbolicToM consistently outperforms both PercepToM and PercepToM+Oracle. We speculate that this performance gap arises because SymbolicToM rephrases the ToM questions into simpler reality questions. For example, the ToM question “Where will Bob look for the celery?” gets rephrased into “Where is the celery?” In contrast, PercepToM addresses the ToM questions as is.
5.4 The Impact of Irrelevant Information
on Perception-to-Belief Inference

### v3

@@ Abstract
Chani Jung  Dongkwan Kim  Jiho Jin  Jiseon Kim Yeon SeonwooYejin ChoiAlice OhHyunwoo Kim KAIST  Amazon  University of Washington  Allen Institute for AI {1016chani, dongkwan.kim, jinjh0123, jiseon_kim}@kaist.ac.kr yseonwoo@amazon.com, yejin@cs.washington.edu, alice.oh@kaist.edu, hyunwook@allenai.org
@@ 1 Introduction
Our contributions are as follows. First, we construct perception-augmented ToM benchmarks which enable the evaluation of the two precursory inferences for ToM in LLMs (§2): perception inference and perception-to-belief inference. Second, using these benchmarks, we show that current LLMs are good at inferring the perceptions of others but struggle to infer beliefs from the perceptual information (§3.2, §3.3, and §4.4). Lastly, we introduce the PercepToM framework to improve LLMs’ ToM reasoning by leveraging their strong perception inference while supplementing their perception-to-belief inference4). Our method enhances LLMs’ performance on ToMi and FANToM, especially on the false belief scenarios (§4.3).
@@ 2.1 Perception Inference and
The precursory inferences for ToM (Rakoczy, 2022) can be understood through the Sally-Anne test, a widely recognized psychological assessment for evaluating ToM (Baron-Cohen et al., 1985). In this test, Sally initially observes a marble in a box but does not witness Anne moving the marble to a basket after she leaves the room.
Current ToM benchmarks predominantly assess LLMs based on their surface-level performance on ToM questions (e.g., “Where will Sally look for the marble when she returns?”), leaving their underlying inference capabilities underexplored. To address this gap, drawing from psychology literature, we refer to the ability to infer others’ perceptions (e.g., “Did Sally see Anne moving marble to the basket?”) as perception inference. Additionally, we define the process of deducing beliefs from perceptual information (e.g., “Sally did not see the marble being moved. Where will she look for the marble when she returns?”) as perception-to-belief inference.
To further investigate these inferences, we construct Percept-ToMi and Percept-FANToM by annotating each character’s perception of information within the context of the two benchmark datasets ToMi (Le et al., 2019) and FANToM (Kim et al., 2023). Annotation examples are presented in Figure 2.
@@ 2.3 Perception-Augmented ToM Benchmarks
To construct Percept-ToMi, we sample 150 story-question pairs for each of the four ToM question types in ToMi [^1] We use the Fixed and Disambiguated ToMi constructed by Sclar et al. (2023), where sentences are inserted to disambiguate the location of containers in the story, and some mislabeled questions are corrected.: first-order true/false beliefs, and second-order true/false beliefs. We automatically annotate perceivers of the scenes in ToMi using SymbolicToM (Sclar et al., 2023) and manually verify the samples. SymbolicToM tracks the witnesses of each scene by maintaining a graphical representation of the true world state, allowing us to obtain the list of perceivers for each scene from its output. However, upon verifying 50 samples of the SymbolicToM output, we identify two types of errors in the perceiver annotations and correct them across our entire dataset. The details of this verification and correction of perceiver annotations are explained in Appendix A.1.
@@
To build Percept-FANToM, we use all of the short conversations in FANToM, but exclude those that cause errors in our perception annotation format. We assume that a character is the perceiver of all utterances that occur from the time they join the conversation until the time they leave. After two of the authors confirmed the criteria for determining the joining and leaving times (Appendix A.2), each data point was manually annotated by one of the authors. They followed the annotation criteria mechanistically, ensuring no subjectivity was involved. Additionally, the authors randomly selected 20 samples to check for any discrepancies in their annotations. The results confirmed that all annotations were consistent between both authors. Based on the annotations of characters’ joining and leaving times, perceivers of each utterance are automatically mapped. The statistics of our perception-augmented ToM benchmarks and the source benchmarks are shown in Appendix B.
@@ 2.4 Task and Evaluation
In order to evaluate the perception inference capability of LLMs, we prompt the models to track characters’ perception of each unit of information in the input context. Specifically, we require the models to respond in the format of a JSON array, which consists of JSON objects containing a unit of information from the context as a key and the perceivers of the information as a value. [^2] We structure the perception inference results in JSON to leverage its parsability and interpretability. Also, recent works use JSON format to improve language model generation quality (Zhou et al., 2023; OpenAI, 2023). We use individual sentences and utterances as the units of information for ToMi and FANToM, respectively. To ensure the generated answers are in the correct format, we provide an example format of the JSON array using a dummy sentence that does not appear in the datasets. The example input prompt is in Appendix C.1.
@@
To evaluate the perception-to-belief inference capability of the models, we provide them with a ground truth perception inference result and then query ToM questions from the original benchmarks. The ground truth perception inference result is provided in the same JSON array format we use to evaluate the perception inference capability of LLMs. The example and detailed explanation of the input prompt can be found in Appendix C.2.
3 Precursory Inferences of ToM in LLMs
3.1 Experimental Setup
We analyze the perception inference and perception-to-belief inference2.1) performances of LLMs on Percept-ToMi and Percept-FANToM (§2.3) with the following metrics and models.
Perception Inference
To evaluate the model-generated perception inference results, we calculate accuracy for a given input context based on the ratio of information units where the model accurately identifies the perceivers. The final perception inference accuracy for a dataset is obtained by averaging the accuracies across all contexts in the dataset. In Percept-ToMi, accuracy is calculated across the stories, with each story paired with a single ToM question. In Percept-FANToM, since multiple questions share a single context, we calculate accuracy across these contexts.
Perception-to-Belief Inference and ToM
We evaluate the perception-to-belief inference and ToM performance of LLMs using the original questions and answers from ToMi (Le et al., 2019) and FANToM (Kim et al., 2023). For ToMi, we measure accuracy by the ratio of correctly answered questions among all story-question pairs. Note that we do not use the joint accuracy proposed in the original ToMi, where a story is counted as correctly answered only if all questions about the story are answered correctly. This is because many of the stories in the Fixed and Disambiguated ToMi (Sclar et al., 2023) do not include all six question types of ToMi. For Percept-FANToM, we report the set:ALL score, which requires the model to correctly answer five types of ToM questions [^3] BeliefQ[Choice], Answerability Q[List], InfoAccess Q[List], Answerability Q[Y/N], InfoAccess Q[Y/N] for the same piece of information within a conversation.
Target Models
3.2 Perception Inference
As shown in Figure 3, most LLMs exhibit high accuracy on perception inference in both Percept-ToMi and Percept-FANToM. The models’ average perception inference accuracy is 0.781 on Percept-ToMi and 0.926 on Percept-FANToM. Also, they exhibit negligible differences in the accuracy between the true belief and false belief scenarios. In ToMi, all models except for GPT-3.5 Turbo and Gemini 1.0 Pro exhibit a gap of less than 0.1 accuracy between the two scenarios. In FANToM, the accuracy gaps between the two scenarios in all models are no greater than 0.014. This result contrasts with the models’ large performance gap in the two scenarios on ToM questions, suggesting that their limited ToM performance in false belief scenarios is not due to the lack of perception inference capability. Detailed results are in Appendix E.
Especially in the ToMi’s true belief scenarios, the two performances exhibit a near-zero correlation (Figure 4). Although moderate correlations appear in other scenarios, the correlation coefficients are not statistically significant. These results imply that LLMs’ perception inference capability is not directly linked to their ToM performance. This contrasts with human adults, where ToM is strictly dependent on perception inference.
3.3 Perception-to-Belief Inference
LLMs struggle with perception-to-belief inference.
Surprisingly, although the ground truth perception information for all characters is provided in this task, models still underperform in false belief scenarios compared to true belief scenarios (see Figure 3). This trend is consistent with their ToM performance. Moreover, their performances on the perception-to-belief inference task are mostly similar to their ToM performances in all scenarios except for the ToMi true belief scenario. The fact that the LLMs hardly benefit from the additional character perception information, which should serve as significant hints for solving ToM questions, suggests that they have limited capability to infer beliefs from perceptions. The exact performances of models are in Appendix E.
This is consistent across all datasets and scenarios (Figure 4). Notably, in FANToM, models exhibit a high correlation between the two performances ($r>0.6$). This correlation likely arises because the two tasks use the same questions. However, since LLMs are showing similar performances in both tasks, we can see that they are not fully leveraging the ground truth perception information in the perception-to-belief inference task.
@@ 4 PercepToM: Grounding ToM Reasoning on Perception
4.1 Framework
According to our experiment results, LLMs perform adequately well in both true and false belief scenarios on perception inference, while they underperform in perception-to-belief inference (§3). Based on these findings, we propose PercepToM, a framework for improving LLM’s ToM reasoning by grounding it in perception information. PecepToM leverages LLM’s strong perception inference capabilities while enhancing its perception-to-belief inference with a simple string-matching rule. PercepToM consists of the following steps as illustrated in Figure 5:
@@ 4.1 Framework
    Perception Inference: The LLM infers which characters perceived each unit of information in the context (e.g., a scene description or an utterance).
@@
    Perspective Context Extraction: Based on the perception inference result from the LLM, PercepToM extracts the perspective context — i.e., the subset of the input context identified by the LLM as perceived by the target character. This process is conducted by a simple string-matching procedure.
@@
If the model correctly performs perception inference, the perspective context will only include what the target character has perceived from the original context – that is, what they believe to be true, based on the principle of rational belief (Baker et al., 2011). When given this isolated context along with the ToM question, the scenario becomes a simple true belief scenario, wherein the LLM has access to the same information as the target character (i.e., information symmetry).
@@
Figure 5: An overview of our PercepToM framework, which enhances LLMs’ ToM reasoning by (1) instructing LLMs to infer the perceivers of each information in the context; (2) aiding their perception-to-belief inference through the perspective context extraction step, which isolates the context perceived by the target character; and (3) allowing LLMs to generate responses to ToM questions based on this perspective context.
@@
SymbolicToM (Sclar et al., 2023) also helps LLM’s ToM reasoning by providing only the context included in the target character’s belief state graph to the model. However, constructing the belief graph in SymbolicToM requires manually crafted algorithms tailored to different types of input. In contrast, PercepToM avoids this requirement by leveraging LLM’s perception inference capabilities, which can handle more diverse and complicated contexts, thereby achieving significantly improved generalizability. The example input and output of each step of the algorithm are provided in Appendix D.
4.2 Experimental Setup
Datasets and Metrics
We evaluate PercepToM and other baseline models on Fixed and Disambiguated ToMi (Le et al., 2019; Sclar et al., 2023) and FANToM (Kim et al., 2023). As evaluation metrics, we use the ratio of correctly answered story-question pairs for ToMi and the set:ALL score for FANToM, as employed in the ToM performance evaluation in the previous section (§3.1).
Baselines
@@ 4.2 Experimental Setup
We compare Vanilla, Chain-of-Thought (CoT; Wei et al., 2022), and System 2 Attention (S2A; Weston and Sukhbaatar, 2023) with PercepToM. Vanilla involves LLM directly answering questions based on the given context, while CoT adds the prompt “Let’s think step by step.” to help the model answer ToM questions. S2A improves the reasoning of LLMs by prompting them to extract only the relevant part of the input context before yielding a final response. By using S2A as a baseline, we compare the effectiveness of the perspective context of PercepToM with the relevant context extracted by LLMs using S2A. We also compare the performance of PercepToM to that of SymbolicToM (Sclar et al., 2023) on ToMi. However, we do not extend this comparison to FANToM, as applying SymbolicToM to input formats other than ToMi is not trivial, given that it is specifically tailored to ToMi’s input format.
Target Models
@@
Since PercepToM leverages the perception reasoning capability of LLMs, we choose models that show reasonable performance on the perception inference task. Specifically, among the eight models, we exclude the bottom two in terms of perception inference accuracy on Percept-FANToM and Percept-ToMi, which are GPT-3.5 Turbo, Claude 3 Haiku, and Gemini 1.0 Pro. As a result, we apply our PercepToM framework to GPT-4 Turbo, GPT-4o, Claude 3 Sonnet, Llama-3 70B Instruct, and Mixtral 8x22B.
4.3 Results
@@
Table 1: PercepToM outperforms the baseline models in most of the scenarios on ToMi and FANToM. Bold indicates the best performance within each language model and scenario (true belief or false belief). Performance comparison between PercepToM and SymbolicToM on ToMi can be found in Appendix F.
The performance of PercepToM is also compared with that of SymbolicToM (Sclar et al., 2023) on ToMi (Appendix F). PercepToM performs comparably to SymbolicToM in false belief scenarios across most LLMs. However, in true belief scenarios, SymbolicToM consistently outperforms both PercepToM and PercepToM+Oracle. We speculate that this performance gap arises because SymbolicToM rephrases the ToM questions into simpler reality questions. For example, the ToM question “Where will Bob look for the celery?” gets rephrased into “Where is the celery?” In contrast, PercepToM addresses the ToM questions as is.
4.4 Impact of Irrelevant Information on Perception-to-Belief Inference